OpenStudy (anonymous):

Tell whether the pair of expressions are equivalent -4+x and x+(-4) Yes or No

OpenStudy (anonymous):

@Adryin

OpenStudy (anonymous):

again the commutative property, this time for addition

OpenStudy (anonymous):

\[a+b=b+a\\ -4+x=x+(-4)\]

OpenStudy (adryin):

These are the same

OpenStudy (anonymous):

Okay! Thanks